Determine the molarity of a NaOH solution when each of the following amounts of acid neutralizes 25.0 mL of the NaOH solution.

5.00 mL of 0.250 M HNO3



1
Expert's answer
2021-03-22T06:45:26-0400

Solution:

The balanced chemical equation:

NaOH + HNO3 → NaNO3 + H2O

According to the equation above: n(NaOH) = n(HNO3)

n = MV


Dilution equation can be used.

Dilution equation is expressed as:

M1V1 = M2V2

where

M = molarity of solution

V = volume of solution


Thus:

M(NaOH) × V(NaOH) = M(HNO3) × V(HNO3)

M(NaOH) × (25.0 mL) = (0.250 M) × (5.00 mL)

M(NaOH) = (0.250 M × 5.00 mL) / (25.0 mL) = 0.05 M

M(NaOH) = 0.05 M


Answer: The molarity of a NaOH solution is 0.05 M.